OpenStudy (anonymous):

Solve by factoring: 15x^2+85x+100=0

OpenStudy (lgbasallote):

divide everything by 5 3x^2 + 17x + 20 = 0 (3x+5)(x+4) = 0 x = -5/3 x = -4

OpenStudy (anonymous):

Danke!